Write the note on spring constant


1
Expert's answer
2021-02-01T08:04:41-0500

Explanation


  • It is the proportionality constant of Hook's law that is described for the regainable deformation of spring and written as "\small F= kx", "\small k: \text{the spring constant}".
  • Dimension can be found as

"\qquad\qquad
\begin{aligned}
\small [k]&=\small \frac{[F]}{[x]}=\frac{MLT^{-2}}{L}=MT^{-2}
\end{aligned}"

  • Unit is commonly written as "\small Nm^{-1}"
  • Stiffness of a spring increases as the string constant increases & vice-versa.
  • It is unique to a given spring & varies between springs with the dimension & the material they are made of.
  • This can be understood by comparing it with the general explanation to the elastic behavior of matter which is written as "\small \sigma=E\epsilon \to \large \frac{F}{A} \small =Y\large\frac{e}{L}\small\to F=\large \frac{AY}{L}\small \implies k=\large \frac{AY}{L}"
  • This form is applicable for all the solid substances within their proportionality limits but the name "spring constant" is usually used for springs only as for a given spring it is constant & noo need of describing the behavior in terms of A, Y or L